FOURIER TRANSFORM INFRARED AND DENSITY FUNCTIONAL THEORY STUDIES OF THE VIBRA-TIONAL FUNDAMENTALS AND STRUCTURES OF Ge_{n} CLUSTERS TRAPPED IN SOLID Ar

\underline{L. DING}, C.M.L. RITTBY, and W.R.M. GRAHAM, Department of Physics and Astronomy, TCU, Fort Worth, TX 76129.

Fourier transform infrared measurements on germanium clusters formed by trapping in solid Ar at 12 K the products from the ablation of Ge rods using a Nd-YAG laser are being carried out in conjunction with density functional theory calculations on Ge_{n} structures, vibrational fundamentals, and isotopic shifts. Overlapping isotopic shifts from the large number of naturally occurring Ge isotopes complicate the FTIR spectra. Positive identification of the Ge_{n} species responsible, assignment of their vibrational fundamentals, and determination of their structures therefore depend on comparison of the observed FTIR spectra with simulated spectra based on DFT predictions. Structures of Ge_{n} clusters are compared with the structures previously determined for their Si_{n} counterparts.